Find the boiling point of a solution containing 45 grams of calcium chloride dissolved in 250 mL of solution.


1
Expert's answer
2021-06-08T21:24:04-0400

"Tb = Kf \u00d7 M"

Kf= 2.53

Moles= 45/111= 0.405moles

M=( 0.405× 250)/1000

M= 0.101M

Tb= 25.3× 0.101

Tb= 25°C